# SADA: Safe and Adaptive Inference with Multiple Black-Box Predictions

### v1

@@ Appendix F Proof of Theorem 2
Appendix G Generating Predictions Using Large Language Models
In this section, we provide step-by-step guidance on generating predictions using large language models, as applied in our empirical studies in Section 5. The input to the LLM consists of four components: a detail section (providing natural language descriptions of each data column), a background (shared across all data points within a dataset), a question (same across all data points), and a list of prompts (10 in total, each reflecting a distinct tone or inquiry style). Consequently, each data point yields 10 separate outputs—one per prompt—and the final prediction is computed by averaging these outputs. Section G.1 and G.2 present example prompts used in the experiments of wine reviews and the politeness evaluation, respectively.
We recommend using at least 80 GB of RAM to run LLaMA-3-70B, though in our experiments, we used Llamma-3-8B as a substitute due to limited resources, which requires approximately 20 GB of RAM. For GPT-4o and DeepSeek-V3, we utilized API calls to improve efficiency and reduce infrastructure requirements.
The approximate data generation speeds for each model in our experiments are as follows: (i) Llama-3-8B: 1000 examples every 10 minutes; (ii) GPT-4o: 1000 examples every 100 minutes; and (iii) DeepSeek-V3: 1000 examples every 15 hours.
G.1 Example prompt for wine reviews
Details: The wine under review is called ”Vermeil 2010 Luvisi Vineyard Zinfandel (Calistoga)”. This wine has a price of \$42.0/bottle of Zinfandel (wine variety) from Calistoga, Napa. It is produced by Vermeil and labeled as Luvisi Vineyard. The wine is described as: Gigantic, huge, enormously extracted Zinfandel. So strong, it’s practically a food group. Offers tiers of raspberries, cherries, blackberries, black currants, raisins, dark chocolate and spices. The alcohol level is very high, giving it a hot finish. It’s a classic example of this old vine, superripe style.
Background: You are a professional wine taster with expertise in U.S. wines. Based on wine descriptions, please evaluate the wine’s quality and assign a score from 80 (lowest score) to 100 (highest score), just like a sommelier would.
Question: Choose an integer between 80 (lowest acceptable quality) and 100 (exceptional quality) to reflect your evaluation. No explanation is needed, and do not give any numbers other than the rating. Your answer must be in JSON format with an integer only, without additional text.
Prompts:
  • Given the wine’s description and background, how would you rate its overall quality?
  • Considering the region, variety, and tasting notes, what score would you assign to this wine?
  • Based on your expertise as a sommelier, where does this wine fall on the 80-100 quality scale?
  • Reflecting on the sensory cues and price point, how would you evaluate this wine?
  • Taking all available details into account, what would be a fair score for this wine?
  • From a wine critic’s perspective, how would you assess the quality of this U.S. wine?
  • Using your tasting intuition and the provided features, what score best represents this wine’s character?
  • Judging from the varietal, region, and tasting profile, what rating would you give?
  • As a seasoned wine evaluator, how would you numerically rate this wine’s craftsmanship?
  • Given the flavor profile and winemaking context, what score reflects your impression of this wine?
G.2 Example prompt for politeness evaluation
Details: The text under review is: ”Doing a redundant ‘atoi‘ after you just did ‘strtol‘ is probably the most twisted insult to the proper use of string-to-int conversion functions in C one can come up with. Why do you see the need to ”re-convert” the value using the broken function ‘atoi‘ when you already have it as ‘val‘ from a proper function ‘strtol‘?” and is from the community: Stack Overflow.Please evaluate its level of politeness based on linguistic features.
Background: This is a post from an online blog. Based on its linguistic features—such as word choice and tone—please evaluate the level of politeness. Then, assign a score from 1(very impolite) to 25(extremely polite), just like a sommelier would.
Question: Choose an integer between 1 (lowest) and 25 (highest) to reflect your evaluation. No explanation is needed, and do not give any numbers other than the rating. Your answer must be in JSON format with an integer only, without additional text.
Prompts:
  • Given the writing style and overall tone of this blog post, how would you rate its level of politeness?
  • Taking into account the language used, the word choices, and the sentiment conveyed, what politeness score would you assign?
  • Imagine you’re an expert in online communication. Based on the phrasing and attitude in this post, what number would you give for politeness?
  • As someone who understands tone and subtext, how polite does this blog post feel to you?
  • Just like a sommelier tasting wine, how would you score the politeness of this post, based solely on its language?
  • If you had to assign a politeness score between 1 and 25 to this writing sample, what would it be?
  • What’s your judgment on the tone of this blog post—how polite does it seem on a scale from 1 to 25?
  • Considering the linguistic subtleties in this post—formality, tone, and word choice—how would you evaluate its politeness?
  • Reflecting on how this post might come across to a casual reader, what politeness score would you give it?
  • Evaluate the post as if you were rating its etiquette. What score feels most appropriate?

### v2

@@ 7 Discussion
Appendix S7 Proof of Theorem 3
Proof.
Theorem 3 follows directly from theorem S4, which provides a non-asymptotic bound for the expected loss. ∎
Theorem S4.
(i) Under Assumption 2, we have that with probability at least $1-\delta$,
$\mathcal{R}(\widehat{{\bm{\theta}}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})\leq\frac{16\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon+\sqrt{\frac{24B^{2}}{n}\log\frac{2}{\delta}}+\frac{2B}{3n}\log\frac{2}{\delta},$
and
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})$- $\mathcal{R}({\bm{\theta}}^{*})\leq\sqrt{\frac{128BC\sqrt{2n}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+8n\gamma B^{2}}{n^{2}}\log\frac{4}{\delta}}$
$+\frac{16\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+\frac{16\sqrt{2(N-n)}C}{N}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{2})}d\epsilon$
$+\sqrt{\frac{24B^{2}(N-n)}{N^{2}}\log\frac{4}{\delta}}+\frac{4B}{3N}\log\frac{4}{\delta},$
where $\gamma=\sup_{{\bm{\theta}}\in\Theta}\operatorname{var}(\widetilde{\ell}_{\bm{\theta}})/\operatorname{var}(\ell_{\bm{\theta}})\leq 1$, and $B$ is the upper bound of the loss function $\ell(f_{\bm{\theta}}(\mathbf{x}),y)$.
(ii) If additionally $\widehat{Y}_{k}=Y$ for some $k\in\{1,\dots,K\}$, we have that with probability at least $1-\delta$,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})$- $\mathcal{R}({\bm{\theta}}^{*})\leq\frac{16\sqrt{2}C}{\sqrt{N}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon+\sqrt{\frac{24B^{2}}{N}\log\frac{2}{\delta}}+\frac{2B}{3N}\log\frac{2}{\delta},$
Proof of Theorem S4.
Recall that the simplified notations
$\ell_{\bm{\theta}}:=\ell(f_{\bm{\theta}}(\mathbf{x}),\mathbf{y}),~~\widehat{\ell}_{\bm{\theta}}:=\frac{N}{N-n}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}),\widehat{\mathbf{y}}),~~\widetilde{\ell}_{\bm{\theta}}:=\ell_{\bm{\theta}}-\frac{N-n}{N}\widehat{\ell}_{\bm{\theta}},$
$\ell_{{\bm{\theta}},i}:=\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),\mathbf{y}_{i}),~~\widehat{\ell}_{{\bm{\theta}},i}:=\frac{N}{N-n}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i}),~~\widetilde{\ell}_{{\bm{\theta}},i}:=\ell_{{\bm{\theta}},i}-\frac{N-n}{N}\widehat{\ell}_{{\bm{\theta}},i}$
We can rewrite the loss functions as
$\widehat{\bm{\theta}}^{\text{nv}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell_{{\bm{\theta}},i},$
and
$\widehat{\bm{\theta}}^{\text{sada}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\widetilde{\ell}_{{\bm{\theta}},i}+\frac{N-n}{N}\frac{1}{N-n}\sum_{i=n+1}^{N}\widehat{\ell}_{{\bm{\theta}},i}.$
Notice that
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})$ $=\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})+\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}}^{*})-\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}}^{*})-\mathcal{R}({\bm{\theta}}^{*})$
$\leq\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})+\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}}^{*})-\widehat{\mathcal{R}}^{\text{nv}}(\widehat{\bm{\theta}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})$
$\leq\sup_{{\bm{\theta}}\in\Theta}\{\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})\}+\sup_{{\bm{\theta}}\in\Theta}\{\mathcal{R}({\bm{\theta}})-\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})\}.$
Denote $g(S):=\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})\right\}$. By the Assumption 2, we know that $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})$ is a sub-Gaussian process with respect to the metric $\frac{Cd_{1}}{\sqrt{n}}$. Applying the classical Dudley’s entropy integral bound, we have
$\mathbb{E}g(S)\leq 8\sqrt{2}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,Cd_{1}/\sqrt{n})}d\epsilon$ $=8\sqrt{2}\int_{0}^{\infty}\sqrt{\log N(\epsilon\sqrt{n}/C,\Theta,d_{1})}d\epsilon$
$=\frac{8\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon.$
Also, by the Assumption 2 and the Talagrand’s concentration inequality (Sen 2018, Theorem 8.7), the following holds with probability at least $1-{\delta}/{2}$,
$g(S)\leq\mathbb{E}[g(S)]+\sqrt{\frac{6B^{2}}{n}\log\frac{2}{\delta}}+\frac{B}{3n}\log\frac{2}{\delta}.$
We can bound $\sup_{{\bm{\theta}}\in\Theta}\{\mathcal{R}({\bm{\theta}})-\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})\}$ in the same way. Combining them together, we know it holds with probability at least $1-\delta$,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})\leq\frac{16\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon+\sqrt{\frac{24B^{2}}{n}\log\frac{2}{\delta}}+\frac{2B}{3n}\log\frac{2}{\delta}.$
We will decompose the SADA estimator $\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})$ into two terms and handle each of them using the same method as above,
$\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})=\frac{1}{n}\sum_{i=1}^{n}\widetilde{\ell}_{{\bm{\theta}},i}+\frac{N-n}{N}\frac{1}{N-n}\sum_{i=n+1}^{N}\widehat{\ell}_{{\bm{\theta}},i}=\widehat{\mathcal{R}}_{3}({\bm{\theta}})+\frac{N-n}{N}\widehat{\mathcal{R}}_{2}({\bm{\theta}}).$
where $\widehat{\mathcal{R}}_{2}({\bm{\theta}})={(N-n)}^{-1}\sum_{i=n+1}^{N}\widehat{\ell}_{{\bm{\theta}},i}$ and $\widehat{\mathcal{R}}_{3}({\bm{\theta}})={n}^{-1}\sum_{i=1}^{n}\widetilde{\ell}_{{\bm{\theta}},i}$. By the same argument as before,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})$ $\leq\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})\right\}+\sup_{{\bm{\theta}}\in\Theta}\left\{\mathcal{R}({\bm{\theta}})-\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})\right\}$
$\leq\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}_{3}({\bm{\theta}})-\mathbb{E}[\widetilde{\ell}_{\bm{\theta}}]\right\}+\sup_{{\bm{\theta}}\in\Theta}\left\{\mathbb{E}[\widetilde{\ell}_{\bm{\theta}}]-\widehat{\mathcal{R}}_{3}({\bm{\theta}})\right\}$
$\quad+\frac{N-n}{N}\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}_{2}({\bm{\theta}})-\mathbb{E}[\widehat{\ell}_{\bm{\theta}}]\right\}+\frac{N-n}{N}\sup_{{\bm{\theta}}\in\Theta}\left\{\mathbb{E}[\widehat{\ell}_{\bm{\theta}}]-\widehat{\mathcal{R}}_{2}({\bm{\theta}})\right\}.$
It follows from the exact same argument as $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})$ that the following holds with probability at least $1-{\delta}/{2}$,
$\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}_{2}({\bm{\theta}})-\mathbb{E}[\widehat{\ell}_{\bm{\theta}}]\right\}+\sup_{{\bm{\theta}}\in\Theta}\left\{\mathbb{E}[\widehat{\ell}_{\bm{\theta}}]-\widehat{\mathcal{R}}_{2}({\bm{\theta}})\right\}$ $\leq\frac{16\sqrt{2}C}{\sqrt{N-n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{2})}d\epsilon$
$\quad+\sqrt{\frac{24B^{2}}{N-n}\log\frac{4}{\delta}}+\frac{2B}{3(N-n)}\log\frac{4}{\delta}.$
Denote $f(S):=\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}_{3}({\bm{\theta}})-\mathbb{E}[\widetilde{\ell}_{\bm{\theta}}]\right\}$. Again, by the Dudley’s entropy integral bound, we have
$\mathbb{E}f(S)\leq 8\sqrt{2}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,Cd_{3}/\sqrt{n})}d\epsilon\leq\frac{8\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon$
By the Talagrand’s inequality, it holds with probability $1-{\delta}/{4}$,
$f(S)\leq\mathbb{E}f(S)+\sqrt{\frac{32BC\sqrt{2n}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+2n\gamma B^{2}}{n^{2}}\log\frac{4}{\delta}}+\frac{B}{3n}\log\frac{4}{\delta}.$
We can bound $\sup_{{\bm{\theta}}\in\Theta}\left\{\mathbb{E}[\widetilde{\ell}_{\bm{\theta}}]-\widehat{\mathcal{R}}_{3}({\bm{\theta}})\right\}$ similarly. Putting all pieces together, it holds with probability at least $1-\delta$,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})$ $\leq\sqrt{\frac{128BC\sqrt{2n}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+8n\gamma B^{2}}{n^{2}}\log\frac{4}{\delta}}$
$\quad+\frac{16\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+\frac{16\sqrt{2(N-n)}C}{N}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{2})}d\epsilon$
$\quad+\sqrt{\frac{24B^{2}(N-n)}{N^{2}}\log\frac{4}{\delta}}+\frac{4B}{3N}\log\frac{4}{\delta}.$
When there exists $k\in\{1,\dots,K\}$ so that $Y=\widehat{Y}_{k}$, we have $\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})=\frac{1}{N}\sum_{i=1}^{N}\ell_{{\bm{\theta}},i}$, which essentially has the same form as $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})$ but replace $n$ with $N$. So following the same logic as before, we have
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})$- $\mathcal{R}({\bm{\theta}}^{*})\leq\frac{16\sqrt{2}C}{\sqrt{N}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon+\sqrt{\frac{24B^{2}}{N}\log\frac{2}{\delta}}+\frac{2B}{3N}\log\frac{2}{\delta}.$
Appendix S8 Additional simulation results
Table S3 and S4 present the bias and coverage probability of the methods examined in the simulation studies described in Section 5.
Table S3: Bias of different methods under varying prediction quality $\gamma$.
Methods Bias
$\gamma=0$ $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.3$ $\gamma=0.4$ $\gamma=0.5$ $\gamma=0.6$ $\gamma=0.7$ $\gamma=0.8$ $\gamma=0.9$ $\gamma=1$
Naive - 0.007 0.007 -0.001 -0.002 -0.001 0.006 -0.004 -0.009 0.002 0.009 -0.003
PPI $\widehat{Y}_{1}$ 0.005 0.009 -0.005 -0.002 -0.001 0.004 0.001 -0.004 0.000 0.002 0.003
$\widehat{Y}_{2}$ -0.001 0.000 -0.003 0.001 -0.003 0.000 -0.002 -0.001 0.005 0.010 0.001
PPI++ $\widehat{Y}_{1}$ 0.007 0.007 -0.001 -0.003 -0.002 0.005 0.000 -0.006 0.001 0.003 0.000
$\widehat{Y}_{2}$ 0.002 0.003 -0.003 0.000 -0.003 0.003 -0.003 -0.006 0.003 0.009 -0.003
SADA both 0.002 0.002 -0.003 -0.000 -0.003 0.003 0.000 -0.005 0.001 0.003 0.000
Table S4: Coverage probability of different methods under varying prediction quality $\gamma$.
Methods Relative efficiency
$\gamma=0$ $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.3$ $\gamma=0.4$ $\gamma=0.5$ $\gamma=0.6$ $\gamma=0.7$ $\gamma=0.8$ $\gamma=0.9$ $\gamma=1$
Naive - .947 .940 .957 .951 .943 .930 .931 .953 .945 .955 .947
PPI $\widehat{Y}_{1}$ .948 .954 .957 .949 .939 .952 .947 .938 .950 .946 .935
$\widehat{Y}_{2}$ .962 .953 .942 .953 .946 .937 .939 .959 .951 .948 .954
PPI++ $\widehat{Y}_{1}$ .942 .939 .955 .954 .931 .935 .936 .941 .942 .958 .935
$\widehat{Y}_{2}$ .955 .944 .943 .945 .944 .929 .924 .950 .938 .951 .945
SADA both .946 .932 .939 .949 .938 .932 .938 .943 .937 .957 .924
Appendix S9 Generating predictions using LLMs
In this section, we provide a step-by-step description of how we generate predictions using the large language models like GPT-4o, Llama-3-8B, and DeepSeek in the empirical study presented in Section 6.1.
The input provided to the LLM consists of four components: a detail section that gives natural language descriptions of each data column, a background section that provides shared task context for all observations for example what kind of text is being evaluated and how politeness is defined, a question section that specifies what the model is asked to predict for every data point and describes the required output format for example a politeness score on a fixed scale., and a set of ten prompts, each designed to represent a different tone or inquiry style. As a result, each data point produces ten separate outputs, one for each prompt, and the final prediction is obtained by averaging these outputs.
Input prompt for generating responses Details: The text under review is: {text}, and is from the community: {source}. Please evaluate its level of politeness based on linguistic features. Background: {backgroud}. Question: {question}. Prompts: {prompt 1},…, {prompt 10}.
Below is an example of the prompts used in the politeness evaluation experiment.
An example prompt used in the politeness evaluation experiment Details: The text under review is: “Doing a redundant ‘atoi’ after you just did ‘strtol’ is probably the most twisted insult to the proper use of string-to-int conversion functions in C one can come up with. Why do you see the need to “re-convert” the value using the broken function ‘atoi’ when you already have it as ‘val’ from a proper function ‘strtol’?”, and is from the community: Stack Overflow. Please evaluate its level of politeness based on linguistic features. Background: This is a post from an online blog. Based on its linguistic features—such as word choice and tone—please evaluate the level of politeness. Then, assign a score from 1 (very impolite) to 25 (extremely polite), just like a sommelier would. Question: Choose an integer between 1 and 25 that reflects your evaluation. No explanation is needed, and do not give any numbers other than the rating. Your answer must be in JSON format with an integer only, without additional text. Prompts: Given the writing style and overall tone of this blog post, how would you rate its level of politeness? Taking into account the language used, the word choices, and the sentiment conveyed, what politeness score would you assign? Imagine you’re an expert in online communication. Based on the phrasing and attitude in this post, what number would you give for politeness? As someone who understands tone and subtext, how polite does this blog post feel to you? Just like a sommelier tasting wine, how would you score the politeness of this post, based solely on its language? If you had to assign a politeness score between 1 and 25 to this writing sample, what would it be? What’s your judgment on the tone of this blog post—how polite does it seem on a scale from 1 to 25? Considering the linguistic subtleties in this post—formality, tone, and word choice—how would you evaluate its politeness? Reflecting on how this post might come across to a casual reader, what politeness score would you give it? Evaluate the post as if you were rating its etiquette. What score feels most appropriate?